40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Mar 21 15:13:44 2017, gautam, Update, IMC, IMC input beam mode matching IMG_7220.JPGIMC_ModeMatch.pdfsingleLensSensitivity.pdfsensitivity.pdfIMCmodeMatch.m
    Reply  Tue Mar 21 21:59:48 2017, gautam, Update, IMC, IMC input beam mode matching IMG_3654.JPGIMG_3655.JPG
       Reply  Wed Mar 22 00:33:00 2017, gautam, Update, IMC, IMC length offset nulling offsetInvestigation.pdfoffset_summing_amp.pdf
          Reply  Wed Mar 22 16:58:25 2017, gautam, Update, IMC, WFS sensing matrix measurements 
             Reply  Thu Mar 23 01:44:53 2017, gautam, Update, IMC, WFS sensing matrix measurements IMC_WFS_segment_TF.pdfIMC_WFS_channels_TF.pdfTFsummary.pdfIMC_WFS_170322.xlsx.zip
                Reply  Thu Mar 23 08:43:11 2017, rana, Update, IMC, WFS sensing matrix measurements 
                   Reply  Thu Mar 23 23:38:58 2017, gautam, Update, IMC, MC SUS damping gains stepped down 
                      Reply  Fri Mar 24 11:26:57 2017, gautam, Update, IMC, MC SUS damping gains restored 
                         Reply  Fri Mar 24 19:04:18 2017, gautam, Update, IMC, Seismic feedforward and WFS MCL_comparison.pdfseis_comparison.pdf
Message ID: 12898     Entry time: Tue Mar 21 21:59:48 2017     In reply to: 12896     Reply to this: 12899
Author: gautam 
Type: Update 
Category: IMC 
Subject: IMC input beam mode matching 

[valera, gautam]

We implemented the plan outlined in the previous elog. The visibility (Pmax-Pmin)/(Pmax+Pmin) calculated with the MC REFL PD levels with the MC locked/unlocked is now ~96% (up from 88%yes). The MC REFL DC level in lock is now ~0.12V (compared to 0.4V). Assuming a modulation depth of 0.1 @ 29.5MHz, about 25% of this (i.e. 0.03V) is from sideband light.

The procedure followed was (see sketch in previous elog for various optic labels):

  1. Move L2 back (towards PMC) by ~2cm.
  2. Walk the beam using M3 and M4 to minimize MCREFL, re-lock IMC, run WFS. 
  3. Move L3 back (towards PMC) by ~2cm.
  4. Repeat steps 2 and 3, the latter with smaller steps, monitor MCREFL DC level.

We could probably tweak the fine positioning of L2 and L3 and improve the efficiency a little more, but the primary objective here was to see if there was any effect on the large common mode offset on the WFS demodulated "SUM" output. Unfortunately, we saw no effect.

Here are two photos of the relevant section of the PSL table before (left) and after (right) our work there:

   

ELOG V3.1.3-